OpenStudy (anonymous):

for the following equation, state the value of the discriminant and then describe the nature of the solutions 17x^2 - 3x + 14 = 0

OpenStudy (bahrom7893):

D = b^2 - 4ac = (-3)^2 - 4*17*14

OpenStudy (bahrom7893):

D = 9 - 952 = -943

OpenStudy (anonymous):

which one of the statements below is correct? 1. the equation has one real solution 2. the equation has two real solutions 3. the equation has two imaginary solutions

OpenStudy (bahrom7893):

x1 = [-b - sqrt(D)]/(2a) x2 = [-b + sqrt(D)]/(2a)

OpenStudy (anonymous):

bahrom thank you so much

OpenStudy (bahrom7893):

np..

OpenStudy (bahrom7893):

x1 = [3 - sqrt(-943)]/(2*17) x2 = [3 + sqrt(-943)]/(2*17)

OpenStudy (bahrom7893):

the answers are imaginary..

OpenStudy (anonymous):

why do you think so?

OpenStudy (bahrom7893):

because of sqrt(-943), you can't take a square root of negative..

OpenStudy (bahrom7893):

3. the equation has two imaginary solutions

OpenStudy (anonymous):

can you solve this one for me please 4sqrtx^12

OpenStudy (bahrom7893):

equals..?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

equals to what?

OpenStudy (anonymous):

that is what we need to find.

OpenStudy (bahrom7893):

[ 4*sqrt(x) ]^12?

OpenStudy (anonymous):

it says use rational exponents to simplify

OpenStudy (bahrom7893):

okay here we go: [ 4*sqrt(x) ]^12 = 4^12 * [ sqrt(x) ]^12

OpenStudy (bahrom7893):

= 4^12 * x^6

OpenStudy (anonymous):

i got the answer as x^3 is that right?

OpenStudy (bahrom7893):

no it's x^6

OpenStudy (bahrom7893):

4^12 * x^6

OpenStudy (anonymous):

i tried it twice but still getting x^3

OpenStudy (bahrom7893):

getting what?

OpenStudy (anonymous):

the answer as

OpenStudy (bahrom7893):

The answer is: 4^12 * x^6 whatever 4 to the 12th power is..

OpenStudy (anonymous):

(2x+6)(9x^2+5x+5) multipy

OpenStudy (bahrom7893):

sorry i was helpin someone else.. someone help out stlouisan..

OpenStudy (anonymous):

thank you Bahrom